# Rationale-Guided Retrieval Augmented Generation for Medical Question Answering

### v1

@@ Abstract
Large language models (LLM) hold significant potential for applications in biomedicine, but they struggle with hallucinations and outdated knowledge. While retrieval-augmented generation (RAG) is generally employed to address these issues, it also has its own set of challenges: (1) LLMs are vulnerable to irrelevant or incorrect context, (2) medical queries are often not well-targeted for helpful information, and (3) retrievers are prone to bias toward the specific source corpus they were trained on. In this study, we present RAG2 (RAtionale-Guided RAG), a new framework for enhancing the reliability of RAG in biomedical contexts. RAG2 incorporates three key innovations: a small filtering model trained on perplexity-based labels of rationales, which selectively augments informative snippets of documents while filtering out distractors; LLM-generated rationales as queries to improve the utility of retrieved snippets; a structure designed to retrieve snippets evenly from a comprehensive set of four biomedical corpora, effectively mitigating retriever bias. Our experiments demonstrate that RAG2 improves the state-of-the-art LLMs of varying sizes, with improvements of up to 6.1%, and it outperforms the previous best medical RAG model by up to 5.6% across three medical question-answering benchmarks. Our code is available at https://github.com/dmis-lab/RAG2
@@
Jiwoong Sohn1  Yein Park1  Chanwoong Yoon1  Sihyeon Park1 Hyeon Hwang1Mujeen Sung2Hyunjae Kim1,†Jaewoo Kang1,3,† 1Korea University  2Kyung Hee University  3AIGEN Sciences {jw_sohn,522yein,cwyoon99,sihyeon-park,hyeon-hwang}@korea.ac.kr mujeensung@khu.ac.kr{hyunjae-kim,kangj}@korea.ac.kr
@@ 1 Introduction
Figure 1: Our RAG2 framework comprises the following three steps. (1) Rationale-based query formulation: the model-generated rationale is used as the query for evidence retrieval instead of the initial query. (2) Balanced retrieval: evidence snippets are retrieved in equal amounts from four corpora: two large ones, PubMed and PMC, and two smaller but crucial corpora, textbooks and clinical guidelines. Subsequently, a reranker is used to rerank the retrieved snippets by cross-encoding the initial query and each snippet. (3) Rationale-guided filtering: a filter model determines whether to incorporating the retrieved snippets into the LLM prompt increases the confidence of the LLM’s rationale (or reduces its perplexity).
However, applying RAG in the biomedical domain presents unique challenges. First, irrelevant or incorrect context can mislead LLMs. Wu et al. (2024). In biomedical texts, the complexity and specialized terminology heighten the risk of retrieving documents that are not relevant or unhelpful. Second, medical queries often struggle to target useful information. Some questions that include extensive patient details, such as medical histories and symptoms, may overwhelm retrieval systems, making it difficult to identify critical diagnostic clues. Conversely, overly brief questions often lack necessary context and depend on implicit medical knowledge, further complicating effective retrieval. Lastly, retriever bias Chen et al. (2021); Dai et al. (2024) also presents a significant challenge, as larger medical corpora often overshadow smaller, specialized ones that may contain critical, up-to-date information. For instance, MedCPT Jin et al. (2023) may show a preference for PubMed documents over clinical guidelines and medical textbooks, and this tendency is further exacerbated by its training on the PubMed corpus.
@@
We evaluate our RAG2 framework using three closed-book medical QA benchmarks—MedQA Jin et al. (2021), MedMCQA Pal et al. (2022), and MMLU-Med Hendrycks et al. (2021)—where no oracle documents are present. Our method significantly enhances the average accuracy of state-of-the-art LLMs of varying sizes, including Llama-3-8B AI@Meta (2024), Meerkat-7B Kim et al. (2024), and GPT-4o OpenAI (2024). Additionally, it consistently improves the performance of baseline LLMs by up to 6.1% across three medical QA datasets and outperforms four recently developed RAG baselines, achieving a notable 5.6% improvement over MedRAG Xiong et al. (2024a) with the Llama-3-8B-Instruct model AI@Meta (2024).
Our contributions are summarized as follows:
@@
    We introduce RAG2, a novel RAG framework for medical QA. Our model incorporates advanced methods in query formulation, retrieval, and filtering processes.
@@
    Our novel rationale-guided filtering method employs a small language model trained with data from rationale perplexity to assess document utility. This approach not only enhances efficiency but also introduces a unique aspect to our methodology.
@@
    In closed-book medical QA benchmarks, we empirically demonstrate that the proposed RAG2 framework significantly outperforms previous methods, including large commercial, small open-source, and medical-specialized models, with an average accuracy improvement of 6.1%, 3.8%, and 0.9%, respectively.
@@ 2.1 Retrieval-Augmented Generation
Basic RAG approaches Lewis et al. (2020); Gao et al. (2023) involve the straightforward addition of retrieved documents to input queries. Advanced methods train LLMs to perform RAG adaptively Wang et al. (2023b); Zhang et al. (2024); Gan et al. (2024). Self-RAG Asai et al. (2024), for instance, trains an LLM with synthetic data from GPT-4 to determine whether to use the retrieved document or not. However, these methods generally require substantial computational resources, making updates challenging.
@@
Several RAG methods leverage the rationales of models to enhance RAG. Speculative RAG Wang et al. (2024b) generates multiple pseudo-answers with rationales and uses a large verifier model to select the best one. RAT Wang et al. (2024a) iteratively performs retrieval and refines the initial rationale to obtain the final answer. However, these approaches come with significant downsides, such as high computational costs due to the use of large models for training and multiple LLM calls, as well as increased latency from the iterative retrieval and decision-making processes.
@@ 3.1 Objective
where $\mathbf{y}_{<t}$ represents the sequence of tokens generated up to time step $t-1$, and $\mathbf{D}^{*}=[\mathbf{d}_{1},\dots,\mathbf{d}_{k}]$ represents the top-k retrieved documents that are most relevant to the query. The retrieval process involves embedding both the query and documents in a shared latent space using a query encoder $f_{\text{q}}$ and an document encoder $f_{\text{d}}$. The most relevant documents are then selected based on their inner product as follows:
@@ 3.2 Rationale-Guided Filtering
Retrieval mechanisms do not always pick the most relevant documents as similarity alone does not ensure a positive contribution to the model’s output. To filter out distractors and obtain only useful document, we train a filtering model to evaluate the retrieved documents before they are presented to the LLM. Our method begins by analyzing the LLM’s outputs when answering questions under two conditions: (1) using its intrinsic parametric knowledge, and (2) using external information retrieved from a corpus. If the model successfully answers a question with the aid of retrieved documents but fails to do so independently, the documents are labeled as “helpful”; otherwise, they are labeled as “not helpful.” This labeling strategy was similarly adopted in previous works Jeong et al. (2024b), but it fails to fully capture the document utility, particularly in cases where the model’s accuracy remains unchanged regardless of whether the document is included or not.
@@ 5.1 Ablation Study
Table 3: Performance of Llama-3-8B-Instruct depending on the filtering method applied. The top-1 snippets are used except for InstructRAG-ICL (top-5)
@@
We investigate how the performance of the baseline model, Llama-3-8B-Instruct, is affected by different filtering methods. To isolate the impact of filtering in this experiment, we fix the top-k value at 1 and used the original query instead of a rationale-based query. For InstructRAG, top-k value is 5 following the default setting. The “GPT-4o” filtering determines whether a retrieved document contributes to answering the question. Our filtering method uses the perplexity of the LLM’s rationale to annotate question-document pairs. As a result, this approach significantly improves Llama’s performance, surpassing Adaptive-RAG by 1.3% on MedMQ and 2.7% on MedMCQA, as shown in Table 3. Additionally, it achieves results similar to GPT-4o while avoiding its API costs.
Effect of Balanced Retrieval
@@
Table 4: Comparison of MedRAG Xiong et al. (2024a) and our balanced retrieval. The top-1 documents are used.
@@ Acknowledgements
This research was supported by (1) the National Research Foundation of Korea (NRF2023R1A2C3004176, RS-2023-00262002), (2) the Ministry of Health & Welfare, Republic of Korea (HR20C0021(3)), and (3) ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(IITP-2024-20200-01819).
@@ A.3 Implementation Details
A.4 Quality of Rationale Queries
Table A2 shows the performance changes based on which model generated the rationale query. Consistently across all backbone models, performance improves most significantly with rationales generated by GPT-4o, followed by Meerkat and Llama-3-Instruct. Although commercial models like GPT-4o provide higher-quality rationales and can further boost performance, relying on them introduces additional dependencies and increases API usage costs.
Rationale Query of
Model Llama-3-8B-Instruct Meerkat-7B GPT-4o
Llama-3-8B 63.4 71.5 73.6
Meerkat-7B 71.3 74.6 78.8
GPT-4o 87.4 88.3 89.8
Table A2: Performance comparison of backbone models using rationale queries generated by Llama-3-8B, Meerkat-7B, and GPT-4o. The results indicate that GPT-4o-generated rationales consistently yield the highest performance improvements across all models, followed by Meerkat-7B and Llama-3-8B.
Figure A2: Performance of Llama-3-8B using different RAG pipelines on ClinicalQA25, which represents a real-world clinical scenario.
A.5 Open-ended Clinical Questions
@@ A.5.1 Metrics
[…]
Figure A3: Comparison of retrieval strategies. We evaluate four approaches: independent corpus retrieval, stacked retrieval (following MedRAG), reranked stacked retrieval, and our proposed balanced retrieval. The balanced approach consistently outperforms others, highlighting the importance of considering corpus-specific properties in medical information retrieval.
@@
$&\text{Precision}(C,R)=\frac{1}{|C|}\sum_{c\in C}\max_{r\in R}\text{cosine\_similarity}(c,r),\\ &\text{Recall}(C,R)=\frac{1}{|R|}\sum_{r\in R}\max_{c\in C}\text{cosine\_similarity}(r,c),\\ &\text{BERTScore}(C,R)=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}.\end{aligned}$
@@ A.5.2 Results
A.6 Balanced Retrieval
We conduct a systematic evaluation of different retrieval strategies to understand their effectiveness in medical QA, with the results presented in Figure A3. First, we analyze each corpus independently to establish baseline performance and assess the utility individual corpus. Next, following MedRAG’s approach, we implement a stacked retrieval method that combines all corpora and selects top documents based on MIPS scores, followed by additional reranking. Finally, we evaluate our balanced retrieval method, both with and without filtering, which is designed to prevent larger corpora from overshadowing smaller but potentially crucial information sources.
Our experiments reveal several important findings: (1) Independent corpus retrieval consistently underperformed, confirming that no single corpus contained sufficient information to optimally answer diverse medical questions. (2) While the stacked approach offers comprehensive coverage by combining all sources, it shows lower performance than our balanced retrieval method. This suggests that simply combining corpora without considering their distinct characteristics and relative strengths can lead to suboptimal results. Even after reranking, the stacked approach does not outperform our balanced retrieval method. (3) Our balanced retrieval approach, particularly when combined with filtering, consistently achieves the best performance across all retrieval scenarios. This demonstrates that deliberately maintaining representation from diverse sources while filtering out irrelevant content is crucial for effective medical information retrieval.

### v2

@@ Abstract
Large language models (LLM) hold significant potential for applications in biomedicine, but they struggle with hallucinations and outdated knowledge. While retrieval-augmented generation (RAG) is generally employed to address these issues, it also has its own set of challenges: (1) LLMs are vulnerable to irrelevant or unhelpful context, (2) medical queries are often not well-targeted for helpful information, and (3) retrievers are prone to bias toward the specific source corpus they were trained on. In this study, we present RAG2 (RAtionale-Guided RAG), a new framework for enhancing the reliability of RAG in biomedical contexts. RAG2 incorporates three key innovations: a small filtering model trained on perplexity-based labels of rationales, which selectively augments informative snippets of documents while filtering out distractors; LLM-generated rationales as queries to improve the utility of retrieved snippets; a structure designed to retrieve snippets evenly from a comprehensive set of four biomedical corpora, effectively mitigating retriever bias. Our experiments demonstrate that RAG2 improves the state-of-the-art LLMs of varying sizes, with improvements of up to 6.1%, and it outperforms the previous best medical RAG model by up to 5.6% across three medical question-answering benchmarks. Our code is available at https://github.com/dmis-lab/RAG2
@@
Jiwoong Sohn1   Yein Park1   Chanwoong Yoon1   Sihyeon Park1 Hyeon Hwang1Mujeen Sung2Hyunjae Kim1,†Jaewoo Kang1,3,† 1Korea University  2Kyung Hee University  3AIGEN Sciences {jw_sohn, hyunjae-kim, kangj}@korea.ac.kr
@@ 1 Introduction
Figure 1: Our RAG2 framework comprises the following three steps. (1) Rationale-based query formulation: the model-generated rationale is used as the query for evidence retrieval instead of the initial query. (2) Balanced retrieval: evidence snippets are retrieved in equal amounts from four corpora: two large corpora, PubMed and PMC, and two smaller but crucial corpora, textbooks and clinical guidelines. Subsequently, a reranker is used to rerank the retrieved snippets by cross-encoding the initial query and each snippet. (3) Rationale-guided filtering: a filtering model determines whether incorporating the retrieved snippets into the LLM prompt can help the model answer correctly and with higher confidence (as indicated by reduced perplexity).
However, applying RAG in the biomedical domain presents unique challenges. First, irrelevant or incorrect context can mislead LLMs. Wu et al. (2024). In biomedical texts, the complexity and specialized terminology heighten the risk of retrieving documents that are not relevant or unhelpful. Second, medical queries often struggle to target useful information. Some questions that include extensive patient details, such as medical histories and symptoms, may overwhelm retrieval systems, making it difficult to identify critical diagnostic clues. Conversely, overly brief questions often lack necessary context and depend on implicit medical knowledge, further complicating effective retrieval. Lastly, retriever bias Chen et al. (2021); Dai et al. (2024) also presents a significant challenge, as larger medical corpora often overshadow smaller, specialized ones that may contain critical, up-to-date information. For instance, MedCPT Jin et al. (2023) may show a preference for PubMed documents over clinical guidelines and medical textbooks, a tendency likely stemming from its training on the PubMed corpus.
@@
We evaluate our RAG2 framework using three closed-book medical QA benchmarks—MedQA Jin et al. (2021), MedMCQA Pal et al. (2022), and MMLU-Med Hendrycks et al. (2021)—where no oracle documents are present. Our method significantly enhances the average accuracy of state-of-the-art LLMs of varying sizes, including Llama-3-8B-Instruct AI@Meta (2024), Meerkat-7B Kim et al. (2024), and GPT-4o OpenAI (2024). Additionally, it consistently improves the performance of baseline LLMs by up to 6.1% across three medical QA datasets and outperforms four recently developed RAG baselines, achieving a notable 5.6% improvement over MedRAG Xiong et al. (2024a) with the Llama-3-8B-Instruct model AI@Meta (2024).
Our contributions are as follows:
@@
    We introduce RAG2, a novel retrieval-augmented generation framework for medical QA. Our method incorporates advanced methods in query formulation, retrieval, and filtering processes that address core limitations in traditional RAG.
@@
    Our rationale-guided filtering method employs a small language model trained with data labeled from differences in rationale perplexity to assess document utility. This approach not only enhances efficiency but also introduces a novel aspect to our methodology.
@@
    Across three medical QA benchmarks, we demonstrate that the proposed RAG2 framework significantly outperforms previous methods, including large commercial, small open-source, and medical-specialized models, with an average accuracy improvement of 6.1%, 3.8%, and 0.9%, respectively.
@@ 2.1 Retrieval-Augmented Generation
Basic RAG approaches Lewis et al. (2020); Gao et al. (2023) involve the straightforward addition of retrieved documents to input queries. Advanced methods train LLMs to perform RAG adaptively Wang et al. (2023b); Zhang et al. (2024); Gan et al. (2024). Self-RAG Asai et al. (2024), for instance, trains an LLM with synthetic data from GPT-4 to determine whether to use the retrieved document or not. However, these methods generally require substantial computational resources and API costs, making model updates challenging.
@@
Several RAG methods leverage the rationales of models to enhance RAG. Speculative RAG Wang et al. (2024b) generates multiple pseudo-answers with rationales and uses a large verifier model to select the best answer. RAT Wang et al. (2024a) iteratively performs retrieval and refines the initial rationale to obtain the final answer. However, these approaches come with significant downsides, such as high computational costs due to the use of large models for training and multiple LLM calls, as well as increased latency from the iterative retrieval and decision-making processes.
@@ 3.1 Objective
where $\mathbf{y}_{<t}$ represents the sequence of tokens generated up to time step t1t-1, and $\mathbf{D}^{*}=[\mathbf{d}_{1},\dots,\mathbf{d}_{k}]$ represents the top-k retrieved documents that are most relevant to the query. The retrieval process involves embedding both the query and documents in a shared latent space using a query encoder $f_{\text{q}}$ and a document encoder $f_{\text{d}}$. The most relevant documents are then selected based on their inner product as follows:
@@ 3.2 Rationale-Guided Filtering
Retrieval mechanisms do not always pick the most helpful documents as similarity alone does not ensure a positive contribution to the model’s output. To filter out distractors and obtain only helpful document, we train a filtering model to evaluate the retrieved documents before they are presented to the LLM. Our method begins by analyzing the LLM’s outputs when answering questions under two conditions: (1) using its intrinsic parametric knowledge, and (2) using external information retrieved from a knowledge base. If the model successfully answers a question with the aid of retrieved documents but fails to do so independently, the documents are labeled as “helpful”; otherwise, they are labeled as “not helpful.” This labeling strategy was similarly adopted in previous works Jeong et al. (2024b), but it fails to fully capture the document utility, particularly in cases where the model’s accuracy remains unchanged regardless of whether the document is included or not.
@@ 5.1 Ablation Study
Table 3: Performance of Llama-3-8B-Instruct depending on the filtering method applied. Top-1 documents are used except for InstructRAG-ICL (top-5)
@@
We evaluate how the performance of the baseline model, Llama-3-8B-Instruct, is affected by different filtering methods. To isolate the impact of filtering in this experiment, we fix the top-k value at 1 and used the original query instead of a rationale-based query. For InstructRAG, top-k value is 5 following the default setting. The “GPT-4o” filtering determines whether a retrieved document contributes to answering the question. Our filtering method uses the perplexity of the LLM’s rationale to annotate question-document pairs. As a result, this approach significantly improves Llama’s performance, surpassing Adaptive-RAG by 1.3% on MedMQ and 2.7% on MedMCQA, as shown in Table 3. Additionally, it achieves results similar to GPT-4o while avoiding its API costs.
@@
Table 4: Comparison with MedRAG Xiong et al. (2024a) and our balanced retrieval. Top-1 documents are used.
Effect of Balanced Retrieval
@@
5.2 Quality of Rationale Queries
Table 5 highlights that rationale queries generated by higher-performing models consistently enhance the accuracy of LLMs on MedQA. Specifically, GPT-4o-generated rationales consistently yield the highest performance improvements across all models, followed by Meerkat-7B and Llama-3-8B. High-quality rationales refine poorly targeted, complex medical questions into precise and informative queries, minimizing the retrieval of non-informative documents and improving overall retrieval effectiveness.
Rationale Query of
Model Llama-3-8B Meerkat-7B GPT-4o
Llama-3-8B 63.4 71.5 73.6
Meerkat-7B 71.3 74.6 78.8
GPT-4o 87.4 88.3 89.8
Table 5: Performance comparison on MedQA of backbone models using rationale queries generated by Llama-3-8B-Instruct, Meerkat-7B, and GPT-4o.
@@ Acknowledgements
This research was supported by (1) the National Research Foundation of Korea (NRF2023R1A2C3004176, RS-2023-00262002), (2) the Ministry of Health & Welfare, Republic of Korea (HR20C0021), and (3) ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(IITP-2025-20200-01819). Mujeen Sung was supported by (4) No. RS-2022-00155911: Artificial Intelligence Convergence Innovation Human Resources Development(Kyung Hee University), (5) No. RS-2024-00509257: Global AI Frontier Lab), and (6) the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program(IITP-2024-RS-2024-00438239).
@@ A.3 Implementation Details
A.4 Open-ended Clinical Questions
@@ A.4.1 Metrics
$\text{ROUGE-L Precision}(C,R)=\frac{LCS(C,R)}{|C|},$
$\text{ROUGE-L Recall}(C,R)=\frac{LCS(C,R)}{|R|},$
$\text{ROUGE-L F1}(C,R)=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}.$
Figure A2: Performance of Llama-3-8B-Instruct using different RAG pipelines on ClinicalQA25, which represents a real-world clinical scenario.
@@
$\begin{aligned} &\text{Precision}(C,R)=\frac{1}{|C|}\sum_{c\in C}\max_{r\in R}\text{cosine\_similarity}(c,r),\\ &\text{Recall}(C,R)=\frac{1}{|R|}\sum_{r\in R}\max_{c\in C}\text{cosine\_similarity}(r,c),\\ &\text{BERTScore}(C,R)=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}.\end{aligned}$
@@ A.4.2 Results
Figure A3: Comparison of retrieval strategies: independent corpus retrieval, stacked retrieval (following MedRAG Xiong et al. (2024a)), reranked stacked retrieval, and our proposed balanced retrieval, with and without rationale-guided filtered retrieval. The balanced approach consistently outperforms others, highlighting the importance of considering corpus-specific properties in medical information retrieval. The highest scores for each retrieval strategy are marked in bold.
A.5 Balanced Retrieval
To validate the effectiveness of balanced retrieval in medical QA, we conduct corpus ablation studies and compare three retrieval strategies: independent corpus retrieval, stacked retrieval following MedRAG Xiong et al. (2024a) with and without reranking, and our balanced method, with and without filtering (Figure A3).
Our experiments reveal several important findings: (1) Independent corpus retrieval consistently underperformed, confirming that no single corpus contained sufficient information to cover diverse medical questions. (2) While the stacked retrieval broadens coverage by combining all sources, it shows lower performance than our balanced retrieval method. This suggests that simply combining corpora can lead to suboptimal results. Even after reranking, the stacked approach does not outperform our balanced retrieval method. (3) Our balanced retrieval, especially when combined with filtering, consistently achieves the best performance across all retrieval scenarios. This demonstrates that deliberately maintaining representation from diverse sources while filtering out unhelpful documents is crucial for effective medical information retrieval.